Mary two student loans total $9000. One loan is at 5% simple interest and the other is at 6% simple interest. At the end of 1 yr, Mary owes $492 in interest. What is the amount of each loan?
Answer by mananth(16949)

Mary two student loans total $9000. One loan is at 5% simple interest and the other is at 6% simple interest. At the end of 1 yr, Mary owes $492 in interest. What is the amount of each loan?
Mary two student loans total $9000.
5% simple interest------------$x
6% simple interest. -----------9000-x
$492 in interest.
..
5x+6(9000-x)=492*100
5x+54000-6x=49200
-x=-54000+49200
-x=-4800
x=$4800====> 5%
balance $4200 at 6%